OpenStudy (anonymous):

There are 8 and 2 over 3pounds of walnuts in a container, which will be divided equally into containers that hold 1 and 1 over 5pounds. This would fill ____ and 4 over 18containers.

OpenStudy (mertsj):

\[8\frac{2}{3}\div 1\frac{1}{5}= \frac{26}{3}\times\frac{5}{6}=\frac{130}{18}=7 \frac{4}{18}\]